OpenStudy (anonymous):

The sides of the cubes are numbered 1 through 6. If they are both tossed, what is the probability that they will both be a 2?

OpenStudy (anonymous):

answers 1/3 1/6 1/36

OpenStudy (anonymous):

1st dice has P(2)=1/6 2nd dice same So 1/6*16=1/36

OpenStudy (anonymous):

So 1/6*1/6=1/36

OpenStudy (anonymous):

ok thanks